OpenStudy (anonymous):

Find the slope for the table of values: (1, 8.5) (2,11) (3,13.5) (4,16) HELP! :)

OpenStudy (saifoo.khan):

\[slope= \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

The choices are: y = 4x + 8.5 y = x + 8.5 y = 2.5x + 6 Y = 4x + 6

OpenStudy (anonymous):

Hey, thanks for all your help. :)

OpenStudy (saifoo.khan):

no problem!

OpenStudy (turingtest):

is anybody else's screen really small?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

No..

OpenStudy (saifoo.khan):

Try zooming in @turing.

OpenStudy (turingtest):

mine shrunk